### v2

# Privacy risk in GeoData: A survey

Abstract.
With the ubiquitous use of location-based services, large-scale individual-level location data has been widely collected through location-awareness devices. The widespread exposure of such location data poses significant privacy risks to users, as it can lead to re-identification, the inference of sensitive information, and even physical threats. In this survey, we analyse different geomasking techniques proposed to protect individuals’ privacy in geodata. We propose a taxonomy to characterise these techniques across various dimensions. We then highlight the shortcomings of current techniques and discuss avenues for future research. Our proposed taxonomy serves as a practical resource for data custodians, offering them a means to navigate the extensive array of existing privacy mechanisms and to identify those that align most effectively with their specific requirements.
De-identified location data, re-identification, geoprivacy, geomasking
doi: ccs: Security and privacy Data anonymization and sanitizationccs: Security and privacy Pseudonymity, anonymity and untraceabilityccs: Information systems Geographic information systemsccs: Information systems Global positioning systemsccs: Information systems Location based servicesccs: General and reference Surveys and overviews
1. Introduction
Data privacy refers to individuals’ protection and control over their personal data, particularly concerning its collection, processing, storage, and sharing by organisations (Houser and Voss, 2018; Rustad and Koenig, 2019). It is crucial for any data that pertains to, is caused by, or initiated by people. Organisations across various business sectors increasingly produce large databases with millions of records, potentially containing detailed and sensitive information about individuals, such as customers, patients, taxpayers, or travellers (Drechsler and Hu, 2021). Consequently, safeguarding the privacy of such data has become a paramount concern.
Privacy violation occurs when personal information is utilised beyond permissible norms. Thus, the concept of data privacy serves as a guiding principle dictating the collection and handling of data based on its sensitivity and significance. However, while contextual integrity elucidates how privacy breaches can manifest within a dataset by analysing adherence to social norms specific to different contexts, it lacks protective mechanisms beyond policies and regulation (Nissenbaum, 2004).
Geospatial data (referred to herein as geodata) is recognised as crucial information across various domains, including epidemiology, medicine, and social science (Fronterrè et al., 2018). At its most detailed level, this data comprises coordinates pinpointing specific locations. For example, geodata can be linked with other information to create geocoded data, which maps individuals’ address details into geographic coordinates. The amount of high-resolution spatial data has been growing enormously due to the integration of positioning capabilities into mobile, wearable, and global positioning systems. With the increasing availability of data about individuals, geodata enables the precise identification of individuals’ locations (Sharad and Danezis, 2013; De Montjoye et al., 2013; Unnikrishnan and Naini, 2013).
As geodata become increasingly accessible at finer resolutions, attacker capabilities in linking geodata to specific individuals or groups continue to grow. Knowing an individual’s location increases the risk of reidentification through reverse geocoding and can seriously violate the individual’s privacy (Brownstein et al., 2006a; Kounadi et al., 2013). For instance, when publishing a data point, coordinates can be simply calculated and connected with an address. Therefore, the address attribute should be treated as confidential information because it can reveal an individual.
The growing need for protecting discrete geographic data results from technological developments and the need to release data online. Hence, numerous studies in the past decade have proposed various approaches to preserve sensitive information, such as geomasking techniques (Zandbergen, 2014). Geomasking techniques are meant to preserve privacy when publishing geodata while maintaining geographic detail to enable precise spatial analysis of the dataset (Kwan et al., 2004a; Gambs et al., 2014; Wang et al., 2022). These techniques aim to obfuscate the real locations of an individual. For instance, by employing geomasking techniques, the address of an individual does not have to be eliminated but can be transferred to a different position, which would result in less accurate data utility but higher privacy.
Despite the various techniques proposed, there are no generally recommended nor approved geomasking techniques. Each presents its unique set of advantages and disadvantages. Furthermore, none of the current masking methods offer a comprehensive solution to protect location privacy. Therefore, a thorough analysis of the capabilities of each technique is imperative to harness them effectively in practical applications. Moreover, ongoing research and refinement are necessary to enhance the efficacy of geomasking strategies in preserving privacy.
Comparison to other works: Several surveys have studied the geomasking techniques. Steffen et al. (Steffen et al., 2023) systematically review the existing literature on data protection strategies for geocoded data and classify them into aggregation, geographic masking, and synthetic data. The survey also discusses various tools for assessing the risk and utility of protected geocoded data. The recent thesis by Redlich (Redlich, 2022) presents a comprehensive evaluation of geomasking methods designed to balance individual privacy protection with the preservation of spatial information utility. This work provides valuable insights into geographic masking methods’ strengths, limitations, and privacy implications, forming a basis for informed decision-making in spatial data analysis. Wang et al. (Wang et al., 2022) conduct a comprehensive evaluation of eight geomasking methods using simulated geospatial data with diverse spatial patterns to assess their effectiveness in privacy protection and analytical accuracy.
Some studies have explored the applicability of geomasking techniques in specific domains. Kamel Boulos et al. (Kamel Boulos et al., 2022) provide a comprehensive overview of the challenges surrounding location privacy and techniques for preserving geoprivacy within the realm of public health interventions and health research that utilise detailed individual geographic data. (Iyer et al., 2023; Broen et al., 2021; Zandbergen, 2014) explore the use of geomasking techniques for perturbing spatial statistics and patient privacy in epidemiologic research and health analytics. Nowbakht et al. (Nowbakht et al., 2022) investigate the challenge of balancing confidentiality and spatial pattern preservation in sharing agricultural data. The thesis by Kekana (Kekana, 2020) focuses on evaluating geomasking methods on presenting the sleeping locations of the homeless without compromising their spatial confidence. Gao et al.(Gao et al., 2019) examine the performance of random and Gaussian perturbation methods for Twitter users sharing their geotagged tweets.
Further, different studies address a specific concern within the geoprivacy domain (Solymosi et al., 2023). Seidl et al. (Seidl et al., 2018) explore the assessment of the risk of false identification in geomasking techniques. Dupre et al. (Dupre et al., 2022) propose to manage privacy risks associated with geocoded data through geospatial disclosure avoidance techniques while Tiwari et al. (Tiwari et al., 2023) investigate the ethical utilisation of geographic information system (GIS) data using geomaking techniques. Kim et al. (Kim et al., 2021) investigate how do different geomasking methods impact perceived disclosure risk and Hu (Hu, 2019) explore the utilisation of Bayesian methods for estimating disclosure risks in synthetic data.
Overall, all the proposed studies discuss the different geomasking techniques and challenges associated with the increasing availability of geocoded data. To the best of our knowledge, there has not yet been a thorough survey of geomasking techniques and classify these techniques under various aspects. This article aims to fill that gap.
Contribution: We present an extensive survey of geomasking techniques. We introduce a taxonomy that reviews geomasking methods across various dimensions, namely privacy techniques, evaluation measures, and practical considerations. At a high level, this taxonomy differentiates between geomasking techniques based on how each technique can anonymise geodata, the diverse evaluation measures applicable to them, and practical aspects. These practical aspects encompass the experimental datasets, purpose-built prototypes designed specifically for implementing these techniques, and the wide range of application domains where these techniques find relevance. The goal of this taxonomy is to facilitate data custodians in navigating the extensive array of existing mechanisms, aiding them in identifying those best suited to their needs. Furthermore, this survey endeavors to highlight any existing gaps in these techniques, offering valuable insights for future researchers to explore and refine. While our survey provides guidance across various aspects, we acknowledge that other taxonomies proposed by different studies could have been equally adopted. Some of these taxonomies may offer more detailed insights for specific inquiries (e.g., scientists interested in exploring knowledge gaps in the field). Additionally, we offer an overview of advanced large language models (LLMs) specifically designed to tackle geospatial data analytics. While these LLMs may not directly address privacy issues, they provide a valuable framework for researchers to explore and integrate privacy considerations, enabling them to address complex geoprivacy issues within their methodologies.
Outline: The rest of this paper is structured as follows. First, in the following section, we provide an overview of geomasking where we discuss different types of geodata and different privacy attacks on geodata. In Section 3 we then present a taxonomy on geomasking where we describe the 10 dimensions we identified that allow us to characterise geomasking techniques. In Section 4 we then survey existing geomasking techniques, and describe how they fit into our taxonomy. We then discuss open issues in existing geomasking techniques and directions for future research in Section 5, and we conclude this paper in Section 6 with a summary of our findings.
2. Background
This section provides necessary background details on geomasking, starting with different types of geodata and then an investigation into different attacks on geodata.
2.1. Different types of Geodata
Geodata comprises information describing objects, events, or features with a location on Earth, recorded alongside a geographic indicator. It primarily exists in two forms: vector data and raster data. Vector data delineates features such as properties, cities, roads, mountains, and bodies of water, represented by points, lines, and polygons. For instance, a visual representation using vector data might include houses represented by points, roads by lines, and entire towns by polygons. Conversely, raster data consists of pixelated or gridded cells, creating complex imagery like photographs and satellite images.
Besides vector and raster data, location information manifests in various other forms, including census data, mobility data, drawn images, and social media data. Census data provides insights into specific geographical areas, facilitating statistical analysis across different locations. Mobility data offers insights into individuals’ locations and movements, which can be collected through tagging locations on social media, using location-tracking apps, interacting with WiFi beacons, or through Bluetooth, Global Positioning System (GPS), or mobile phone records. This data is valuable for conducting geospatial analysis. Drawn images, such as Computer-aided design (CAD) images, provide information about buildings or other structures, offering geographical insights. Furthermore, social media data can include information about people’s movements as they post about their visited places.
2.2. Different Privacy attacks on Geodata
Geodata privacy refers to a set of principles and practices put in place to protect the privacy of individuals, groups, communities or organisations whose data is represented in geographically explicit formats (Kamel Boulos et al., 2022). With geodata increasingly available at finer resolutions and attackers continuously advancing their capabilities in linking geodata to specific individuals or groups, the risk of a privacy breach escalates, especially when precise geographic locations are known, particularly in areas with low population density. The emergence of various geodata sources introduces new concerns for privacy breaches, as traditional disclosure avoidance mechanisms may overlook spatial characteristics (Bertino et al., 2008; Solymosi et al., 2023). In this section, we explore a range of privacy attacks/breaches targeting geodata.
Figure 1. Fundamental process of address geocoding and reverse address geocoding.
2.2.1. Reverse Geocoding
Geocoding is the process of assigning geographic coordinates to addresses or place names, enabling spatial data analysis and visualisation. However, releasing geographic information at the individual level can pose significant privacy risks. For instance, disclosing a street address can allow individuals to be easily identified through directories and property databases. Similarly, publishing coordinates enables pinpoint locations on maps, potentially revealing sensitive information. Whether in print or digital format, maps containing such data can inadvertently disclose addresses and compromise privacy.
Reverse geocoding occurs when the address details for a published location are determined. These address details serve as powerful keys for linking the data associated with those coordinates to specific and identifiable information from other sources. For example, in the context of census data, the respondents’ locations can be inferred and reverse geocoded if a census area has cells with low values, potentially leading to inadvertent disclosure. Reverse geocoding can facilitate the re-identification of individuals, as it allows the conversion of geographic coordinates into street addresses. Once a street address is obtained, it becomes possible to use common directories to associate that address with one or several individuals. Figure 1 illustrates an example of geocoding and reverse geocoding processes for addresses (Kounadi et al., 2013).
Several research works have shown the re-identification of individual addresses using reverse geocoding. In 2006, Brownstein and others (Brownstein et al., 2006a) were able to correctly identify more than 70% of the addresses in a published map using manual reverse geocoding techniques in GIS. The same authors showed similar accuracy in identifying addresses can be achieved using semi-automated reverse geocoding based on image analysis (Brownstein et al., 2006b). In realistic examples, a group of researchers from the US correctly identified most of the original residents of mortality locations of Hurricane Katrina using a published map in a local newspaper (Curtis et al., 2006). Further, another study of crime incidents in Vienna, Austria, showed that individuals could be accurately re-identified using online reverse geocoding and online address and telephone directories (Kounadi et al., 2013).
2.2.2. Geocoding of administratively-masked data
Helderop et al. (Helderop et al., 2023) propose a geocoding method for unmasking administratively-masked data where specific identifying features, like addresses, are obscured to protect privacy. The proposed approach utilises a medoid-based technique, aiming to minimise spatial uncertainty associated with the masking process. The method utilises a master address point dataset to generate a list of candidate addresses for each masked observation, and a minimum bounding polygon is drawn around these candidates. The medoid, representing the central-most point, is then identified within this polygon and assigned as the geocoded address. The approach outperforms commercial geocoding software, providing higher accuracy and a spatial confidence metric for each result.
2.2.3. Inference attack
An inference attack corresponds to a process by which an adversary that has access to some data related to individuals (and potentially some auxiliary information) tries to deduce new personal information that was not explicitly present in the original data. For geographical data, the inference attacks have been studied for a long time (Sharad and Danezis, 2013; De Montjoye et al., 2013; Unnikrishnan and Naini, 2013). The most common inference attack type applied on geodata is de-anonymising attacks (Gambs et al., 2014). The de-anonymising attack aims to re-identify individuals using their anonymised location data.
Gambs et al. (Gambs et al., 2014) introduce a de-anonymising attack strategy, where an adversary tries to infer the identity of a specific individual through analysis of their mobility traces. More precisely, this attack method demonstrates a high success rate in re-identifying individuals whose movements are documented in an anonymous dataset. This success is contingent upon the adversary having access to mobility traces of the same individuals, observed during a training phase, which they can employ as background information.
Wang et al. (Wang et al., 2016) propose a new model to profile users’ spatiotemporal mobility behaviours and use the model to launch the de-anonymisation attack. The authors use the User Hidden Markov Model to profile the user movements and used an algorithm to rank the users based on the probability that each individual matches with a given profile.
A recent work by Eshun and Palmieri (Eshun and Palmieri, 2022) proposes two de-anonymisation techniques using Hidden Markov Models. In their attacks, both the temporal and spatial influences on user mobility trajectories are considered for predicting user movements. These predictions are then used to create a profile for each individual, and the re-identification of individuals based on these user profiles is achieved through a divergence measure. An experimental evaluation on real-world mobility traces shows that their proposed attack can achieve over 80% of accuracy in re-identifying individuals.
2.2.4. Differencing attack
A differencing attack is possible with variables for which there are multiple different plausible coding schemes, where the categories in those coding schemes are not nested but instead overlap. For example, given a table with information on 20- to 25-year-olds and a table with information on 20- to 24-year-olds, the difference between the two tables will reveal information about 25-year-olds only. If the difference between cells contains a small number of individuals, then a disclosure is more likely.
Differencing attacks occur when data is released according to a geospatial hierarchy such that data in different granularity levels can be combined to reconstruct data at a finer granularity or identify the location of an observation. This situation may occur for maps with different geographical coding potentially allowing more information to be revealed about individuals in the overlaps than intended from each individual table. Although it could happen with tabular data, the issue most commonly comes up with geodata due to the publicly available geospatial information. The result of this differencing is that whilst a single map may be considered safe in isolation, this may not be the case for multiple maps when overlain with one another.
3. A Taxonomy of Geomasking
In this section, we outline our proposed taxonomy encompassing three primary dimensions of geomasking techniques: privacy techniques, evaluation measures, and practical aspects, as illustrated in Fig. 2. As previously discussed in Section 1, our aim in developing this taxonomy is to thoroughly analyse existing geomasking techniques, the measures to evaluate their performance, potential datasets for experimental evaluations, existing purpose-built prototypes of the techniques, the diverse domains where these techniques can be applied, and to identify gaps in existing geomasking techniques. Such analysis will assist data custodians in navigating existing methods and prototypes to identify the most suitable ones based on their specific requirements. In the following subsections, we discuss each dimension in detail.
Figure 2. Taxonomy of geomasking.
3.1. Privacy techniques for Geomasking
Geomasking is a class of methods for changing the geographic location of an individual in an unpredictable way to protect confidentiality, while trying to preserve the relationship between geocoded locations and occurrence events. In this section, we analyse different privacy techniques proposed for geomasking. Following earlier studies (Armstrong et al., 1999; Stinchcomb, 2004; Zandbergen, 2014; Seidl et al., 2018; Kekana, 2020; Broen et al., 2021; Wang et al., 2022; Redlich, 2022; Kamel Boulos et al., 2022; Iyer et al., 2023), we group geomasking techniques into five major categories based on how each technique is treating the data points to provide privacy: (1) Random perturbation, (2) Masking with pre-defined locations, (3) Affine transformation, (4) Aggregation and clustering, and (5) Synthetic data generation. Next, we explore different techniques that are proposed under each of these categories and discuss their potential and limitations. The existing literature for each category is further discussed in Section 4.
3.1.1. Random Perturbation
Random perturbation moves individual points from their original location to a random distance in a random direction (Allshouse et al., 2010). When the generated masked points may be close to the original points and for a small proportion of the geomasked points, the adversary can re-identify individuals by reversing geocode. Next, we explain several techniques that can be categorised under random perturbation.
3.1.1.1 Density-adaptive displacement
In this category, the displacement distance is inversely proportional to the underlying population density. In the geomasking literature, the population densities can be considered either as uniform or heterogeneous.
Figure 3. Circular masking with fixed radius (left), Circular masking with random radius (middle), and donut masking (right).
Uniform population density distribution: The methods under this category assume the population density is distributed homogeneously. Figure 3 illustrates the different methods under this category.
  1. (1)
    Circular masking with a fixed radius: This method generates random locations on a circle around the original points by a predetermined displacement distance while ensuring that the perturbed location points do not fall within the circle (Kwan et al., 2004a).
  2. (2)
    Circular masking with a random Radius: This method randomly displaces points both in direction and distance within a circular area from their original positions, ensuring they remain confined within the boundaries of the circle (Armstrong et al., 1999; Kwan et al., 2004a). All locations inside the circle have an equal chance of being selected.
  3. (3)
    Donut Masking (random perturbation within an annulus): The donut method extends current methods of random displacement by ensuring that an address is not randomly assigned on or too near its original location. In donut masking, each geocoded address is relocated in a random direction by at least a minimum distance, but less than a maximum distance. In population-based donut masking, each point is moved at a distance inversely proportional to the underlying population density. This means that points within regions of lower population density are displaced over larger distances compared to those situated in areas of higher population density. Essentially, the displacement magnitude is dynamically adjusted based on the concentration of the population, ensuring that privacy protection is maintained while minimising spatial error. Consequently, points in sparsely populated regions undergo more substantial repositioning compared to points in densely populated urban areas, thereby striking a balance between data privacy and spatial accuracy (Stinchcomb, 2004; Allshouse et al., 2010).
Heterogeneous population density distribution: The methods under this category assume the population density is distributed non-homogeneously.
  1. (1)
    Spatially adaptive random perturbation (SARP): This technique considers the spatial variation of the population at risk when displacing points by random distances and in random directions (Lu et al., 2012). It reduces unnecessary noise in high-dense population zones by using relatively large and small population zones for low and high population densities, respectively. The SARP geomasking technique defines the perturbation zone based on the actual distribution of residential addresses (risk location) instead of people (risk population). It employs ”donut-shaped” spatially varied perturbation zones rather than ”pancake-shaped” which allows a point to be moved to a nearby location within its immediate surroundings. Utilising actual street addresses the SARP geomasking technique could offer better privacy protection compared to geomasking based on population size.
  2. (2)
    Weighted random perturbation (WRP): This technique has the potential to preserve spatial distributions by displacing points shorter distances in more populated areas (Kwan et al., 2004b). The weighted random mask aimlessly rotates every data point on a circle, applying the variable radius $r$, where $r$ depends on the underlying population density. This approach imposes a maximum displacement limit based on the distance to each point’s $k$-th nearest neighbour.
Figure 4. Gaussian displacement (left), Bimodal Gaussian displacement (middle), and location swapping (right).
3.1.1.2 Gaussian deviation for displacement magnitude
There are several methods that can be considered under this category.
  1. (1)
    Gaussian displacement: This method moves location points to a random direction, and the magnitude of displacement conforms to a Gaussian distribution (Cassa et al., 2008; Fanshawe and Diggle, 2011; Zandbergen, 2014). However, the Gaussian displacement’s main challenge lies in the limited point displacement due to its concentration around the mean. Figure 4 (left plot) illustrates this method.
  2. (2)
    Bimodal Gaussian displacement: This technique is an adaptation of Gaussian displacement, employing a bimodal Gaussian distribution to determine random distances for the masked locations (Cassa et al., 2006; Zandbergen, 2014). It shares similarities with donut masking but offers a probability distribution that is less uniform. Figure 4 (middle plot) illustrates this method.
  3. (3)
    Density–based Gaussian spatial skew: In this method, the location points are adjusted by applying a random offset derived from a Gaussian distribution, where the standard deviations of the distribution are inversely related to the population density in the local area (Cassa et al., 2006). Utilising local demographic data allows the proposed anonymisation system to shift individuals in heavily populated regions by shorter distances compared to those in sparse areas. Consequently, addresses can undergo minimal distortion while upholding a designated level of k-anonymity.
3.1.1.3 Multiple risk factor adaptive displacement
This category incorporates various factors in calculating displacement distances, including but not limited to population density. Triangular Displacement is a method that can be considered under this category. The triangular displacement method considers population density for displacement distances and incorporates multiple factors such as individual re-identification risk, data sensitivity levels, research types, the presence of quasi-identifiers, and potential exposure to external sources to calculate displacement distances (Murad et al., 2014). However, this technique has limitations. It may be vulnerable to edge effects and could result in masked locations falling into impractical or unrealistic areas, thereby impacting data confidentiality. Despite these limitations, this method represents a step forward in geomasking techniques, offering a more nuanced and context-aware approach to safeguarding spatial data in research and analysis.
3.1.2. Masking with pre-defined locations
The proposed methods under this category move a given point to a new location, which is randomly chosen from all the possible locations of a similar type that exist within a specific region (i.e., a circle or an annulus) surrounding the original location. Location swapping methods offer advantages over random methods by providing a more realistic approach to selecting displaced locations, ensuring these locations fall within predefined areas of interest rather than arbitrary placements (i.e., bodies of water or uninhabited are). This enhances anonymity and preserves spatial patterns more akin to unmasked locations. However, if a malicious actor can ascertain the buffer distance used for relocation, there remains a risk of re-identifying the original locations (Zhang et al., 2017).
  1. (1)
    Location Swapping: This method swaps a given location/point to a new location based on the similar geographic characteristics of a point within a specified neighbourhood (Zhang et al., 2017). The method first predefined a radius around the location to be swapped. Then, the method randomly selects a location among all the possible locations that fall within the defined radius. This method is similar to the random perturbation within a circle for a couple of reasons: 1) Within the predefined radius, all the locations are considered with equal probability, making the swapping to any location possible within the defined radius, and 2) the radius can be defined based on the underlying population density distributions. However, the location swapping method considered only the locations that have predefined locations (such as residential addresses) as possible candidates for swapping, and the radii can be varied based on the local population densities. Figure 4 (right plot) illustrates an example of location swapping.
  2. (2)
    Location Swapping within Donut: This method represents another variation of location swapping, employing the concept of donut masking to set a minimum distance, preventing points from being swapped with any locations within this specified range (Zhang et al., 2017). The determination of this minimum distance can be based on local population densities. The authors employed an internal buffer size half that of the external buffer.
  3. (3)
    Street Masking: Similar to Location swapping, in the street masking method, the original location is replaced with coordinates of a point sampled from a set of points within a predefined radius (Swanlund et al., 2020b). Unlike location swapping, which relies on residential addresses, this method utilises road networks from a given map for sampling. During point selection, any points not in proximity to a drivable road (such as in rural areas) are further adjusted by moving them to the nearest applicable road. Similar to location swapping, street masking can leverage population densities to determine the radius, assuming areas with high population density also exhibit high road density and vice versa. Nonetheless, in areas with high road density, this method may experience longer runtimes due to the larger number of points to be masked.
Street masking offers a convenient solution for geographic masking by leveraging road network data, which is widely available globally. By aligning data points with street networks, street masking preserves topological continuity between original and masked locations and minimises false attribution risks, ensuring realistic shifts without arbitrary displacements. However, its reliance on road networks may introduce biases towards urban and built environments, potentially skewing analytical outcomes. In contrast, location swapping offers a viable alternative by allowing for more controlled displacement of data points without the inherent biases associated with street-centric masking. Location swapping is preferable in scenarios where avoiding point displacement onto roads is essential. This method overcomes issues related to urban biases and road network dependencies by swapping points with verified neighbors or alternative locations. Street masking lacks intuitive control over the search depth parameter, leading to varying masking distances influenced by local road configurations and density. This lack of control complicates the process of determining optimal settings, posing a challenge for users seeking consistent results across different datasets and study areas.
Figure 5. Affine transformation methods, Flipping (left), rotation (middle), and scaling (right).
3.1.3. Affine-Transformation
The transformation techniques, which are also known as isomasks techniques, apply basic geometric transformations to move spatial point data (Ajayakumar et al., 2019). Here, it can be differentiated between rotate, scale and translate, stochastic mechanism, and concentration of isomasks. Affine transformations change the original location by either twisting every point by a set angle (referred to as rotation), moving every point by a set increment (referred to as translation) or moving every point by scaling constant (referred to as scale). Next, we describe these methods in more detail, and Fig. 5 illustrates some of these methods.
  1. (1)
    Translation: This method adds a random constant displacement value to both x and y coordinates while ensuring the distance between coordinates remains the same (Wang et al., 2022). All points will be shifted to a fixed distance and direction.
  2. (2)
    Flipping: This method manipulates the orientation of axes, both on a local and global scale. There are several flipping techniques, such as global horizontal flipping, which flips the map’s horizontal axis; global vertical flipping, which flips the map’s vertical axis; global axes flipping, which flips both horizontal and vertical axes; and local random flipping, which randomly alters the orientation of horizontal, vertical, or both axes within grid cells to hide location points (Gupta and Rao, 2020; Leitner and Curtis, 2004, 2006).
  3. (3)
    Rotating: This method rotates each point by a fixed angle ($0\leq\theta\leq 360$), centering the rotation around either the origin of the coordinate system or an arbitrarily chosen point (Wang et al., 2022).
  4. (4)
    Scaling (Resizing): This method multiplies a constant value with both x and y coordinates, which remains consistent for both axes (Wang et al., 2022). Consequently, this process alters the distances between the points and all points will be expanded by a scaling factor.
  5. (5)
    Horizontal shear: This method employs a linear transformation to horizontally shear the location points (Armstrong et al., 1999; Broen et al., 2021). Each point is shifted along its x-axis, oriented at a 45° angle relative to its initial position concerning the center of the point distribution.
Affine transformations preserve the spatial structure of the data, proving useful for spatial analysis and visualisation methods such as clustering, while maintaining the number of records and avoiding random relocation of data (Ajayakumar et al., 2019). However, in these techniques, if the transformation matrix is known, retrieving the coordinates of the original records from the perturbed dataset is simple. Therefore, affine transformations are not a widely used methods (Wang et al., 2022).
3.1.4. Spatial aggregation & Clustering
Spatial aggregation and clustering techniques are commonly used to protect individual privacy by reducing spatial resolution, making it difficult to pinpoint exact locations and preventing the disclosure of precise geographical information. However, these techniques can also result in a loss of data granularity, which may restrict detailed analysis and reduce overall location accuracy.
Spatial aggregation aggregates individual location points within specific geographic units or areas, such as cities, zipcodes, or administrative areas, to ensure anonymity and prevent the disclosure of exact locations. This technique operates by decreasing the spatial resolution of the data, thereby modifying the original data patterns. The goal is to generate generalised records that consolidate information within defined territories, protecting the identities of specific locations. However, this process may compromise the utility of published records as the detailed patterns found in high-resolution data might not be as discernible in the aggregated information. Figure 6 illustrates different spatial aggregation techniques.
Clustering aims to reveal inherent patterns within spatial data by grouping similar data points based on geographical proximity or shared characteristics. Clustering maintains the overall characteristics of an area without pinpointing precise locations, preserving the general patterns while ensuring individual data privacy. However, due to the computations applied directly on original data, clustering could expose sensitive information especially in datasets containing personal or confidential data. This can lead to unintended re-identification of individuals within the dataset, posing risks of privacy breaches and discrimination. Adversaries can exploit clustering results to deduce individuals’ group memberships or glean details about the data distribution, even without direct access to the data (Canbay and Sever, 2015).
Figure 6. Grid center masking (left), quadtree grid (middle), and voronoi masking (right) techniques.
3.1.4.1 Grid-based aggregations
The techniques in this category involve the realignment of individual data points onto a grid system (Leitner and Curtis, 2004). It typically involves relocating each original data point to specific positions within designated grid cells. These relocations may take place at either the centroid or vertices of the grid cells, each being aligned with a defined side length. The efficacy of methods within this category is notably influenced by two crucial parameters: (1) the size of the aggregation units and (2) the boundaries utilised to define these units.
In grid-based aggregations, the grid cell size could be fixed or varied. Grid line masking and grid center masking are two popular methods that use fixed-size grid cells. The grid line masking method relocates location points to the nearest edge of the grid cell in which they are contained (Gupta and Rao, 2020). The grid center masking method relocates the location points to the centroid of the cell in which they are presently situated (Leitner and Curtis, 2004).
  1. (1)
    Quadtree hierarchical geographic data structures is a method that uses varying-size grid cells. This method employs a quadtree methodology on an initial 1km2 European grid and population data points to create a grid with varying sizes that can be adjusted according to local population density(Lagonigro et al., 2017). In this approach, densely populated areas are divided into smaller square areas to ensure accurate data dissemination, while sparsely populated areas are grouped into larger areas to prevent the identification of individual data.
  2. (2)
    Military Grid Reference System (MGRS) protects sensitive point-based geospatial data with a focus on preventing the matching of points with their true underlying observations, such as street addresses. This method involves converting coordinates to the Military Grid Reference System (MGRS) and using digit switching for masking, allowing encryption at different spatial precision levels (Clarke, 2016). The selection of the digit switching combination minimises differences in aggregate descriptive spatial statistics between the masked and unmasked data, with emphasis on the nearest neighbour statistic. Despite the large runtime, this method protect sensitive geospatial data, overcoming the disadvantages of geometric point masking methods, such as the lack of control over parameters, compromised spatial analysis of distribution, and susceptibility to violations of geoprivacy associated with the use of high-resolution imagery and accurate data from Global Navigation Satellite Systems (GNSS). The proposed digit-switching method, which operates effectively across multiple nested spatial scales, addresses these limitations by allowing the users to choose a distribution that closely mimics the spatial distribution of unmasked data, ensuring suitability for spatial analysis based on aggregate descriptive statistics.
3.1.4.2 Spatial tessellation-based techniques
Spatial tessellation methods encompass techniques for dividing geographic areas into smaller, more manageable units while retaining the spatial integrity of the original data.
  1. (1)
    Voronoi masking: Voronoi polygons define areas where the boundaries are equidistant between surrounding points or the interior of the polygons is closer to the corresponding point than to any other points (Seidl et al., 2015). In Voronoi masking, each point is snapped to the closest part along the edges of its corresponding Voronoi polygon. An advantage of this technique is that where the density of original points is higher, the points are moved a shorter distance, resulting in patterns that more closely resemble the original data. The Voronoi masking pattern is strongly linked to the distribution of points within the specified study area. Another advantage of Voronoi masking is that some points in adjacent polygons will be snapped to the same location, which can increase their k-anonymity.
  2. (2)
    Adaptive areal elimination (AAE): This method establishes zones with a minimum level of K-anonymity (Kounadi and Leitner, 2016). Initially, data points are grouped into polygons if available. Subsequently, a disclosure value for ”RoRi” (risk of re-identification) is determined, representing the minimum K-anonymity threshold at which sensitive information can be revealed. The method involves a dissolving process, where polygons with RoRi values below the set threshold are merged with neighbouring polygons until all polygons reach or exceed the threshold. This merging process follows a spatial guideline based on shared borders, primarily prioritising neighbours with the longest shared border. In the case of equilateral polygons, they undergo the dissolving process with all adjacent polygons. This process results in the creation of K-anonymised areas where the disclosure threshold is met. Location points within these areas can be relocated to the centroids or randomly distributed within the K-anonymised zones.
  3. (3)
    Adaptive areal masking (AAM): This method represents an enhanced iteration of AAE (Charleux and Schofield, 2020). In AAM, specific anonymisation polygons are constructed for individual points by iteratively merging them with the nearest centroid-based polygon, followed by the next closest, until the k-anonymity threshold is met. This approach preserves data more effectively than AAE because aggregation stops when the desired k-anonymity level is reached. Additionally, it minimises point displacement by initially merging smaller polygons. Although merged polygons may not always be adjacent, this method is more resource-efficient than verifying adjacency. It further conserves resources by excluding polygons with no points to anonymise, which can be a majority in some applications. Lastly, the process is amenable to parallelisation as each point is treated individually, enhancing efficiency.
  4. (4)
    Adaptive voronoi masking (AVM): This method integrates the principles of AAE and VM to preserve data utility, considering topological polygon relationships and underlying topography, and achieve a high level of spatial k-anonymity to mitigate re-identification risks(Polzin, 2020). A key strength of AAE lies in accounting for underlying population density, while VM excels in displacing the Original Data Points (ODP), generating masked patterns similar to the ODP’s distribution. However, both VM and AAE exhibit drawbacks: they overlook the underlying topography, potentially placing data points in illogical locations such as lakes or forests, leading to false re-identification risks. AVM bridges these gaps by integrating the benefits of AAE’s population density consideration and VM’s ODP displacement method. It moves ODPs to the closest segments of their corresponding Voronoi polygon within the merged AAE-polygon, mitigating the risk of displacing points to areas with different population thresholds. Additionally, AVM addresses topographical concerns by relocating data points to the nearest street intersection, minimising false re-identification and illogical placements.
  5. (5)
    RDV masking: This method, known as the three-layer RDV masking model, is based on rudimentary triangulation and Voronoi generation(Gupta and Rao, 2020). The masking process is centralised, applying masking to specified nodes at one point. The key advantage of this solution is its ability to not only efficiently mask points but also retain the spatial pattern of the original data, considering various indicators beyond population density. The RDV masking method is structured into three distinctive layers, each contributing to the overall masking process. The first layer, termed layer R, initiates by spatially distributing a given set of points throughout the region, ensuring the preservation of the original point configuration for later comparison and analysis. Moving to layer D, the process progresses to computing the Delaunay triangulation of the given point set, utilising a Divide and Conquer strategy to efficiently accomplish this phase, albeit with a specific computation duration. Finally, layer V, the last phase, focuses on separating individual points to their nearest segment along the edges of the corresponding Voronoi polygon. This step is pivotal in upholding the original density patterns of the points, employing Fortune’s Algorithm for computation, which also entails its specific computation duration. Together, these layers encapsulate a comprehensive approach to masking, offering spatial distribution, triangulation, and precise segmentation to maintain original data patterns effectively.
  6. (6)
    Zip4codes aggregation: This method uses Zip4 (postal) codes as a spatial aggregation unit for sharing fine-scale health data while addressing privacy concerns (Ajayakumar et al., 2023). The Zip4 aggregation preserves the underlying spatial structure, making it potentially suitable for geospatial analysis.
3.1.4.3 Differentially private clustering
Differential privacy (DP), recognised as the current state-of-the-art provable privacy technique for mitigating the risk of identity disclosure by incorporating randomisation and noise addition (Dwork, 2006). By leveraging randomisation techniques coupled with noise addition, DP significantly diminishes the disclosure risk associated with these methods, offering robust protection for sensitive data. An algorithm is said to be differentially private if the algorithm produces essentially the same answers regardless of the presence or absence of an individual in the dataset.
Figure 7. K-means clustering (left) and differentialy private k-means clustering (right).
DP has recently been used for querying geospatial datasets in a privacy-preserving manner (Xiao and Xiong, 2015; Harris, 2020). In addition to its application in statistical queries, DP has been employed in geospatial clustering. Clustering geospatial data can inadvertently expose vulnerabilities. For example, the disclosure of cluster centroids during the iterative clustering process can lead to the potential leakage of information about individual data points, particularly when adversaries have background knowledge or auxiliary information enabling them to conduct inference attacks (Lu and Shen, 2020).
DP techniques could inject noise into the clustering process such that the centroids may move slightly or significantly from their original positions depending on the magnitude and distribution of the injected noise (Yan et al., 2022; Wang et al., 2017; Qardaji et al., 2013; Lu and Shen, 2020). Figure 7 illustrates both the original k-means clustering (on the left) and the differential privacy-based k-means clustering (on the right) applied to the same dataset. To mitigate such risks and safeguard individual privacy, considerable research effort has been dedicated to integrating differential privacy into clustering algorithms.
The main drawback of using DP algorithms for data analysis is that, like other perturbation methods, they do not preserve the accuracy of the algorithms applied to the original data. In fact, they tend to deviate more from the results based on the original data in proportion to the privacy guarantees they provide (Balcan et al., 2017). Further, DP implicitly assumes that all information about an individual participant is contained within a record and that records are independent. Thus, if the dataset contains multiple location records per person the the DP mechanism needs to be design to capture all these records of an individual. If location records are correlated, the privacy protections may weaken, as the DP mechanism may inadvertently leak information due to the interdependence of records.
3.1.5. Synthetic data generation
Synthetic datasets aim to preserve the spatial characteristics and patterns present in the original data while preserving the individual’s privacy. Synthetic data can be generated in a number of ways (Bauer et al., 2024). Some synthetic data generation techniques include (1) Generative AI; these techniques are based on Generative Pre-trained Transformers (GPT), Generative Adversarial Networks (GANs), or Variational Auto-Encoders (VAEs) which learn the underlying distribution of real data to generate similarly distributed synthetic data (Zhao et al., 2021; Wu and Biljecki, 2022). (2) A rules engine that can create synthetic data via user-defined business rules. Intelligence can be added to the generated data by referencing the relationships between the data elements, to ensure the relational integrity of the generated data (McKenzie et al., 2023). (3) Entity cloning is another technique that extracts data from the source systems of a single entity (e.g., patient) and masks it for compliance. It then clones the entity, generating different identifiers for each clone to ensure uniqueness. (4) Data masking is also a possible option that replaces personally identifiable information with fictitious, yet structurally consistent, values, such as addresses or coordinates (Nowok and Dibben, 2018). The objective of data masking is to ensure that sensitive data cannot be linked to individuals, while retaining the overall relationships and statistical characteristics of the data.
Some other techniques, such as Copula models, could discover correlations and dependencies within the datasets, and then use these correlations and dependencies to generate new, realistic data. Data augmentation (such as flipping, rotation, scaling, and translation) could also be applied as supplementary techniques to existing values to create new data. Finally, Techniques such as random sampling and noise injection, which add data points from known distributions, could also be used to create new data points that closely resemble real-world data (Sakshaug and Raghunathan, 2014).
3.2. Evaluation measure for Geomasking
The performance of the geomasking techniques can be evaluated in terms of how efficient and effective they are when applied to geospatial datasets. The efficiency of a geomasking technique evaluates how scalable a masking technique is on a large dataset with millions of records, while the effectiveness of a geomasking technique is measured by the accuracy and the protection it can provide for an individual. Next, we discuss different measures that have been proposed to evaluate geomasking techniques in terms of scalability, accuracy, and privacy.
3.2.1. Scalability
Scalability can be evaluated using measures that are dependent on the computing platform and networking infrastructure used, or measures that are based on the number of data points analysed (Redlich, 2022). Common measures include run time, memory space, and communication size.
Apart from these measures, the scalability can also be evaluated theoretically using computational complexity measures. This includes the computation and communication complexities (if multiple data providers are available) that measure the overall computational efforts and cost of communication required in the geomasking process. Generally, the big O-notation is used to specify the computation complexity (Papadimitriou, 2003). Given $n$ is the number of records in a database, the big O-notation of represents $O(log\ n)$ logarithmic complexity, $O(n)$ linear complexity, $O(n\cdot log\ n)$ log-linear complexity, $O(n^{2})$ quadratic complexity, $O(n^{c})$ polynomial complexity, $O(polylog\ n)$ polynomial logarithmic complexity, and exponential complexity, where $c>1$.
3.2.2. Accuracy
Table 1 presents our classification of metrics for evaluating the accuracy of geomasking techniques. These metrics are categorised into three groups: Spatial Point Pattern & Distribution Analysis, Displacement Distance Analysis, and Spatial Autocorrelation Analysis.
Table 1. Metrics for evaluating the accuracy of geomasking techniques
Accuracy Metric Category Metric
Spatial Point Pattern & Distribution Analysis Average Nearest Neighbor (ANN)
Kernel Density Estimation (KDE)
Point density estimation (PDE)
Ripley’s K
L-function
Cross-K function
Minimum convex polygon (MCP)
Standard deviation ellipse (SDE)
Global divergence index (GDi)
Local divergence index (LDi)
Spatial cluster detection
Spatial central tendency and dispersion (normal distribution)
Spatial central tendency and dispersion (non-normal distribution)
Displacement Distance Analysis Displacement distance
Displacement statistical summary
Spatial Autocorrelation Analysis Global Moran’s I
Local Moran’s I
3.2.2.1 Spatial Point Pattern & Distribution Analysis
This category assesses points’ spatial arrangement and distribution in both the original and masked datasets.
  1. (1)
    Average Nearest Neighbour (ANN): ANN is a spatial analysis method used to quantify the arrangement of points within a dataset(Iyer et al., 2023; Wang et al., 2022; Ajayakumar et al., 2023; Broen et al., 2021; Seidl et al., 2015; Zandbergen, 2014). It calculates the average distance of each point to its nearest neighbour, offering valuable insights into the spatial pattern of the data. This metric is particularly useful for identifying clustering at a global level, providing valuable information about the spatial distribution and clustering tendencies of the dataset.
  2. (2)
    Kernel Density Estimation (KDE): KDE is a robust point pattern analysis technique designed to create smooth density surfaces from point feature datasets(Iyer et al., 2023; Wang et al., 2022; Seidl et al., 2018; Ajayakumar et al., 2023; Zandbergen, 2014). It is particularly useful for detecting clusters and evaluating the impact of displacement, as seen in geomasking. KDE allows for the quantitative assessment of spatial clustering. This comparison not only aids in identifying location tendencies but also provides an estimation of the extent to which the original spatial patterns are preserved after masking.
  3. (3)
    Point density estimation (PDE): PDE entails constructing density surfaces or maps from point feature datasets to visually depict spatial patterns (Wang et al., 2022).
  4. (4)
    Ripley’s K: This metric analyses spatial point process data, revealing clustering patterns at various local distances (Dixon, 2001; Ajayakumar et al., 2023; Swanlund et al., 2020b). It summarises point patterns across multiple scales, allowing the determination of clustering occurrences at different distances and underlying spatial patterns.
  5. (5)
    L-function: This metric is a normalised version of Ripley’s K that evaluates the expected count of points within a multi-dimensional sphere with a radius $r$, normalised by the volume of the sphere (Dixon, 2001; Broen et al., 2021). It determines whether the observed distribution of points in close proximity to a particular location indicates either clustering or dispersion, beyond what would be anticipated under random chance.
  6. (6)
    Cross-K function: This metric assesses the spatial clustering patterns between two distinct point patterns (Kwan et al., 2004a; Seidl et al., 2015; Zandbergen, 2014). The Cross-K function quantifies the expected number of points from the reference pattern ($j$) within a specified distance of an arbitrary point in the target pattern ($j^{\prime}$), normalised by the overall density of points within the reference pattern ($j$). This measure explores relationships among multiple point patterns in spatial statistics.
  7. (7)
    Minimum convex polygon (MCP): MCP is a spatial analysis method used to define the minimum convex hull that encompasses a set of points or features within a dataset, presenting the spatial arrangement of features within a dataset (Wang et al., 2022).
  8. (8)
    Standard deviation ellipse (SDE): SDE is utilised to visually represent the spatial distribution of features in a dataset (Yuill, 1971; Wang et al., 2022; Seidl et al., 2018). It serves as a tool for measuring the spread or variability of a point pattern and assessing the concentration of geographic coordinates.
  9. (9)
    Global divergence index (GDi): GDi is specifically designed for summarising disparities between spatial data (Kounadi and Leitner, 2015; Seidl et al., 2018; Nowbakht et al., 2022). GDi quantifies the dissimilarity of masked data in relation to the original data, characterising the spatial information divergence by considering three essential components: the divergence of spatial mean, the orientation, and the major axis of the first standard deviational ellipse. GDi produces an average divergence index that effectively represents the dissimilarity of the obfuscated point pattern to its original counterpart.
  10. (10)
    Local divergence index (LDi): LDi is based on the symmetrical difference between nearest neighbor hierarchical clustering and Getis-Ord Gi* hot spots in comparison to the original point distribution, providing a means for concise point distribution comparisons (Kounadi and Leitner, 2015; Seidl et al., 2018; Nowbakht et al., 2022). It calculates divergence by evaluating the disparity in hotspot areas between the obfuscated and original points.
  11. (11)
    Spatial cluster detection: This metric evaluates the presence of clusters within a set of spatial data points (Seidl et al., 2018; Ajayakumar et al., 2023). Commonly used spatial clustering tools include the Cuzick-Edwards statistic (Cuzick and Edwards, 1990), SaTScan circular clustering detection (Kulldorff, 2006), and GeoMEDD(Curtis et al., 2020).
  12. (12)
    Spatial central tendency and dispersion (normal distribution): Metrics such as minimum (min), maximum (max), mean, mode, median, standard deviation (SD), and Standard Deviational Ellipse are fundamental statistical measures offer insights into the impact of geomasking methods on spatial distortion. Calculated for both the original and masked datasets, these measures enable a comprehensive understanding of how the central tendency and variability of spatial data are influenced (Redlich, 2022; Nowbakht et al., 2022).
  13. (13)
    Spatial central tendency and dispersion (non-normal distribution): Measures such as min, max, median, 1st quartile, 3rd quartile, and the interquartile range (IQR) are useful for characterising the central tendency and dispersion of point data within a dataset (Nowbakht et al., 2022). These metrics allow for a detailed assessment of distortion in datasets with non-normal distributions, contributing significantly to refining geomasking methodologies and interpreting spatial data in various contexts.
3.2.2.2 Displacement Distance Analysis
This category measures the distance of displacement between original and masked locations.
  1. (1)
    Displacement distance: This metric is employed to measure the distance between the masked location and its original location (Swanlund et al., 2020b; Redlich, 2022). It provides insights into the extent of spatial distortion introduced by the masking process.
  2. (2)
    Displacement statistical summary: This metric is utilised to quantify the number or percentage of points within a dataset exhibiting distinct displacements (Nowbakht et al., 2022).
3.2.2.3 Spatial Autocorrelation Analysis
This category examines the spatial correlation of both original and masked data points.
  1. (1)
    Global Moran’s I: This metric is used to assess global-level spatial autocorrelation, providing insights into the distribution pattern of a variable across geographic space (Moran, 1950; Broen et al., 2021; Redlich, 2022; Zandbergen, 2014). Ranging from -1 (indicating complete separation) to 1 (reflecting complete clustering), this index serves as a valuable tool for quantifying the degree of spatial clustering or dispersion in a dataset.
  2. (2)
    Local Moran’s I: This metric is used to assess the presence of local spatial autocorrelation within a geographic dataset (Anselin, 1995; Broen et al., 2021). In contrast to the Global Moran’s I, Local Moran’s I offers a finer-grained analysis by examining the similarity between a specific spatial unit and its neighbouring areas. This metric is particularly valuable for uncovering patterns of spatial autocorrelation that may be concentrated in specific regions rather than uniformly spread across the entire geographic area. By identifying hotspots and cold spots, Local Moran’s I aids in the identification of localised trends, clusters, or outliers, contributing to a more nuanced understanding of spatial relationships within the dataset.
3.2.3. Privacy
In this section, we discuss several metrics that can be used to measure the privacy of the geomasking techniques.
  1. (1)
    Spatial k-anonymity: Spatial k-anonymity is a crucial metric in assessing the disclosure risk of geomasked datasets (Allshouse et al., 2010; Seidl et al., 2015; Iyer et al., 2023; Wang et al., 2022; Seidl et al., 2018; Dupre et al., 2022; Broen et al., 2021; Redlich, 2022; Ajayakumar et al., 2023; Zandbergen, 2014). This method evaluates the likelihood of identifying a specific record within a given spatial region by calculating the number of potential geospatial records in proximity to that record. Spatial k-anonymity leverages reverse geocoding as a disclosure threat mechanism.
    One approach to quantifying spatial k-anonymity entails examining the number or percentage of obfuscated points within a specified distance, considering varying values of $k$. In the geomasking context, the anonymisation region is created as a buffer with a radius ($r$) centred at the masked location, where $r$ corresponds to the distance between the original and masked location. A higher k value reduces the probability of successfully re-identifying the original location, indicative of enhanced geomasking performance. Spatial k-anonymity can also be determined by counting the number of location points within a radius ($r$) for each specific point, where $r$ represents the maximum displacement distance resulting from the masking approach. This calculation is applicable when considering the weight of displacement in masking and evaluating the level of privacy achieved after masking.
    Spatial k-anonymity is often determined using the $n$-th nearest neighbour calculation, referring to the number of potential locations closer to the masked location than to the original location. This approach essentially reveals the empirical estimate of spatial k-anonymity, providing insights into the reidentification risk associated with masked datasets. Additionally, spatial k-anonymity can be measured as the average distance to k nearest neighbors for the masked point. A lower distance to k neighbors suggests a lower probability of association with the original location, indicating a higher level of privacy.
  2. (2)
    Topological identification risk: Apart from spatial k-anonymity, this measure considers the topological information in masked data (Seidl et al., 2018). The metric assesses the privacy risk in geomasking through the identification of topological relations between masked points. This measure conceptualises false identification risk considering factors such as topological relations, the existence of a single nearest neighbour location, and the density of neighbouring locations. The identification risk could be categorised into various scenarios, including points interior or disjoint to correct or incorrect locations, on the boundary of correct and incorrect locations, and disjoint to locations with correct or incorrect nearest neighbours.
Apart from the above measures, Redlich (Redlich, 2022) highlights additional re-identification risk assessments. These measures include linking the geodata with a dataset containing longitudinal information, mean of multiple releases, and reversing masking techniques. Linking geodata with longitudinal information involves mapping masked geospatial records with a dataset containing direct identifiers about individuals, such as names and addresses.
For the linking process, two attack schemes have been proposed. (1) Graph theoretic linkage attack (Kroll, 2015) uses graph theory to link masked and original coordinates, framing the problem as a maximum clique issue. (2) Graph matching attack (Vidanage et al., 2020) uses graph matching mechanisms to identify matching records in encoded and plaintext datasets. The attack utilises features like node-based, edge-based, and structural features, applying cosine locality-sensitive hashing for identifying matching records. The second measure, mean of multiple releases, involves using various dataset versions to determine the original coordinates’ location. The idea is based on the central limit theorem, which states that as the number of observations increases, the average of the masked coordinates will fall closer to the original point. The third measure, reversing masking techniques, aims to reverse a masking method, such as affine transformation, by testing all possible parameters to find the original location.
3.3. Practical aspects of Geomasking
The final three dimensions cover practical aspects of geomasking techniques, including the datasets used for experimental evaluations, how a solution was implemented, and various application domains where geomasking can play a pivotal role.
3.3.1. Datasets
To evaluate the practical aspects of geomasking techniques with regard to their expected performance in real-world applications, evaluations should ideally be done on databases that exhibit real-world properties and error characteristics. However, in some occasions, due to the difficulties of obtaining real-world data that contain geospatial information, synthetically generated geo databases could also be used. Table LABEL:tab:datasets provides details of several datasets that have been used by different studies for experiments with geomasking techniques. It is important to note that though these datasets are available publicly, some of these datasets might have ethical obligations.
Table 2. Datasets used by different studies for experiments on geomasking techniques.
Dataset Description Repository Link Used by
European data The official portal for European data https://data.europa.eu/data/datasets/?locale=en (Polzin, 2020)
French public data Open platform for French public data. (Houfaf-Khoufaf et al., 2021) utilised below datasets - COVID-19 virological test results on
https://www.data.gouv.fr/fr/datasets/old-donnees-relatives-aux-resultats-des-tests-virologiques-covid-19/
- Base National Address on https://www.data.gouv.fr/fr/datasets/base-adresse-nationale/
https://www.data.gouv.fr/fr/datasets/ (Houfaf-Khoufaf et al., 2021)
US Census Bureau data This repository provides data related to People, Places, and Economy; covering over 100,000 different geographies: States, counties, places, tribal areas, zip codes, and congressional districts in the United States. For each, they cover topics including, but not limited to, education, employment, health, and housing. - (Wang et al., 2022) utilised United States Census Bureau, 2021 - (Dupre et al., 2022) utilised United States Census Bureau, 2019 (Income and Poverty in the United States: 2019) https://data.census.gov/ (Wang et al., 2022; Dupre et al., 2022)
US GIS data The NYS GIS Clearinghouse is an evolving searchable repository of GIS data and mapping resources. (Zhang et al., 2017) utilised actual residential datasets from three U.S. counties in their analysis, specifically examining data from Jackson County in Oregon, Travis County in Texas, and Wake County in North Carolina. https://data.gis.ny.gov/ (Zhang et al., 2017)
U.S. Census Bureau’s geographic spatial data This repository represents the U.S. Census Bureau’s geographic spatial data, which has been developed using Topologically Integrated Geographic Encoding and Referencing system (TIGER) geospatial data as the primary source. https://www.census.gov/programs-surveys/geography/guidance/tiger-data-products-guide.html (Su et al., 2017)
US GIS data The NYS GIS Clearinghouse is an evolving searchable repository of GIS data and mapping resources. (Zhang et al., 2017) utilised actual residential datasets from three U.S. counties in their analysis, specifically examining data from Jackson County in Oregon, Travis County in Texas, and Wake County in North Carolina. https://data.gis.ny.gov/ (Zhang et al., 2017)
U.S. Department of Transportation (USDOT) the National Address Database (NAD), established by the USDOT and its partners, provides addresses aggregated from state, local, and tribal governments. These addresses are compiled to provide accurate and up-to-date information critical for various purposes, including transportation safety, Next Generation 9-1-1 services, mail delivery, permitting, and school siting. https://www.transportation.gov/gis/national-address-database (Lin, 2023)
Australian Government open data Data.gov.au serves as the primary repository for open government data in Australia. Individuals can retrieve anonymised public data released by federal, state, and local government agencies. https://data.gov.au/search (Redlich, 2022)
Geoscape Geocoded National Address File (G-NAF) Geoscape G-NAF is a trusted source of geocoded address database for both Australian businesses and governmental entities, comprising more than 50 million submitted addresses condensed into 15.4 million G-NAF addresses. It is built and maintained by Geoscape Australia using independently examined and validated government data. (Redlich, 2022) utilised August 2019 - PSMA Geocoded National Address File (G-NAF) from https://data.gov.au/dataset/ds-dga-e1a365fc-52f5-4798-8f0c-ed1d33d43b6d/distribution/dist-dga-32be073c-338b-40df-9842-7cd91a25d960/details?q= https://data.gov.au/dataset/ds-dga-19432f89-dc3a-4ef3-b943-5326ef1dbecc/details (Redlich, 2022)
Earth Engine Data Catalogue The public data catalogue of Earth Engine encompasses a range of standard raster datasets related to Earth science. (Clarke, 2016) utilised a public GNSS track from a hike on the Tennessee Valley trail in the Golden Gate National Recreation Area, California, retrieved from the user-contributed data for Google Earth. The track consists of 360 points specified in geographic decimal degrees on the WGS84 coordinate system. https://developers.google.com/earth-engine/datasets/catalog (Clarke, 2016)
Deaths recorded in Lawrence, Massachusetts from 1911 to 1913 This repository, provided by (Broen et al., 2021), contains geocoded data on all deaths that occurred in Lawrence, MA, from 1911 to 1913. Additionally, it provides the code for applying various spatial perturbations and generating descriptive spatial statistics. https://github.com/broenk/Spatial_Perturbation (Broen et al., 2021)
GIS data for the city of Lawrence, Massachusetts This repository provides GIS data for the city of Lawrence, Massachusetts. https://data.mass.gov/browse/gis (Broen et al., 2021)
GIS data for Orange County, North Carolina This repository provides GIS data for Orange County, North Carolina. (Allshouse et al., 2010) employed the E911 database for Orange County, North Carolina (Orange County GIS Division 2007), which includes census block groups exhibiting diverse population densities. https://www.orangecountync.gov/1610/Land-Records-GIS-Services (Allshouse et al., 2010)
Bike sharing dataset (NY, USA) This repository provides citi bike trip data, including details on their riding locations, preferred timing, ride distances, popular stations, and the days of the week with the highest ride frequency. https://citibikenyc.com/system-data (Hasan et al., 2017)
Austrian statistical data The platform data.statistik.gv.at provides various datasets sourced from the official Austrian statistical data dictionary. These datasets adhere to the Principles Of Open Data, ensuring their accessibility in machine-readable formats. In the study by (Seidl et al., 2015), energy consumption data for each household in the Hermagor district of Carinthia in southern Austria was utilised. This dataset encompasses 1945 residential records, represented by the centroids of georeferenced buildings supplied by the respective communities in the district. https://www.statistik.at/datenbanken/statcube-statistische-datenbank (Seidl et al., 2015)
Open platform for research data (OSF) OSF serves as a freely accessible and open platform designed to facilitate research and foster collaboration by providing access to projects, data, and materials that could be beneficial to the research process. - Synthetic clusters, addresses, and masked data provided by (Swanlund et al., 2020b): https://osf.io/6uqkv: - Simulated data provided by (Wang et al., 2022): https://osf.io/hxd5w/ https://osf.io/search?resourceType=File (Swanlund et al., 2020b; Wang et al., 2022)
Figshare Figshare is an open cloud-based platform used for managing, sharing, and discovering research outputs, including data, figures, code, presentations, and other types of scholarly content. - A synthetic address-level dataset provided by (Lin, 2023): https://figshare.com/articles/dataset/Geo-indistinguishable_masking_Enhancing_privacy_protection_in_spatial_point_mapping/23632443 https://osf.io/search?resourceType=File (Lin, 2023)
3.3.2. Implementations
This dimension specifies the applications and prototypes that have been developed and implemented for various geoprivacy techniques in order to conduct experimental evaluation. Some solutions proposed in the literature provide only algorithmic details, but they have not been evaluated experimentally, or no details about their implementation have been published. Table 3 provides a brief description of existing tools implementing different geomasking techniques.
Table 3. Different geomasking tools that have been implemented to ensure geoprivacy.
Geoprivacy Tool Description Repository link
GeoPriv (Pablo Duque and Pugliese, 2023) A QGIS 3+ [^1] https://qgis.org/en/site/ plugin that offers three location privacy methods (i.e., Spatial Clustering (Zurbaran and Wightman, 2017), NRandK (Zurbarán et al., 2018), and Laplace Noise (or Geo indistinguishability) (Andrés et al., 2013)). These methods are mainly designed for location-based services and focus on Vector Layers with Point geometries, utilising obfuscation and k-anonymity. https://github.com/Diuke/GeoPriv
MaskMy.XYZ (Swanlund et al., 2020a) A web application designed for easy geographic masking. This tool uses donut geomasking and includes metrics for privacy protection and information loss, enabling users to adjust masking parameters iteratively based on these metrics. https://maskmy.xyz/
PrivyTo (McKenzie et al., 2022) This tool is designed to protect privacy when sharing personal location data. Users can generate multiple obfuscated views using various geomasking techniques (i.e., random perturbations), tailoring each for specific recipients. The application utilises encryption for secure data transfer. https://privyto.me/site
Privy (Ajayakumar et al., 2019) A geomasking tool mainly designed for healthcare. It transfers key location information to a different geographic location, analysing the data, and finally returns the results to the original location. Privy relocates coordinates using a combination of random translations and rotations while maintaining the connectivity of spatial point data. https://github.com/ghhlab/confidentiality
MapSafe (Sharma et al., 2023) A web-based application to share encrypted geospatial information that uses donut masking and hexagonal binning for dataset obfuscation, followed by a multi-level encryption scheme. This tool enables data owners to share data at their preferred level of detail. https://www.mapsafe.xyz/
StppSim (Adepeju, 2023) A tool designed for synthesising fine-grained spatiotemporal crime records. This tool allows to generate synthetic data from a sample dataset or based on user-provided spatiotemporal characteristics. https://github.com/cran/stppSim
GeoMasker (Chen et al., 2017) A geospatial tool designed to balance the protection of geoprivacy and the accuracy of spatial patterns within a GIS. The tool employs carefully calibrated geomasking parameters, including grid size (GS), estimated K-anonymity (K), and D statistics, to mask the residential locations within GIS software. GeoMasker is valuable for health analytics, enabling secure information processing and sharing while preserving privacy and spatial patterns. http://idv.sinica.edu.tw/tachien/geomasker
3.3.3. Application Domains
Geospatial information, spanning a diverse range of location-based data, plays a pivotal role across various application domains, including but not limited to disaster management, health analytics, urban planning and management, and agriculture. These distinct sectors are explored in greater detail in the subsequent subsections.
3.3.3.1 Disaster Management
Geospatial technologies play a pivotal role in disaster management, offering a multifaceted toolkit that aids in various aspects of the disaster lifecycle. Geospatial technologies significantly enhance decision-making, resource allocation, and coordination across all stages of disaster management (Ghosh, 2023). Early warning systems are facilitated through the integration of real-time data, such as weather patterns and seismic activities, while crowdsourced data from mobile applications and social media supplement geospatial information, providing valuable real-time insights during disasters. The interoperability of geospatial tools supports collaboration among various agencies, and simulation tools aid in training and capacity building for disaster management personnel. Additionally, the integration of geospatial technologies with IoT devices enhances real-time monitoring of critical infrastructure, contributing to the overall preparedness and resilience of communities in the face of disasters (Lwin et al., 2019).
The utilisation of geospatial technologies, such as remote sensing and satellite imagery, poses a risk of directly or indirectly observing private property and capturing sensitive personal information without the possibility of informed consent. In situations where visual data is collected and disseminated, individuals may be exposed to privacy violations, placing them in vulnerable positions. Moreover, the absence of robust privacy and data protection standards, especially in the context of data sharing, heightens the potential misuse of geospatial data. This ethical dilemma underscores the importance of addressing privacy issues to ensure responsible and transparent practices in the use of GIS for emergency management (Kicior, 2023; Sutedi et al., 2021).
3.3.3.2 Health Analytics
Geospatial technology plays a crucial role in healthcare by providing spatially explicit information about environmental exposures, lifestyle patterns, and disease risks. Over the last three decades, advancements in Earth observations, spatial data quality, and software usability have enabled the integration of geospatial tools in various health studies (Faruque, 2022; Bhoda, 2023; Shah, 2023). This technology not only contributes to a comprehensive understanding of the complex interplay between genes, lifestyle, and environment but also extends its impact on crucial aspects of public health. Geospatial data serves as a powerful tool for disease surveillance, allowing the collection and analysis of data to monitor disease outbreaks and inform timely responses (Faruque, 2022). It aids in identifying at-risk populations, guiding the allocation of resources during outbreaks (Shah, 2023).
Furthermore, geospatial technology enhances health promotion efforts by offering detailed information on the accessibility of healthcare services, helping policymakers make informed decisions about facility placement to ensure equitable access for all (Sahana et al., 2022). In addition, it plays a key role in evaluating and improving healthcare delivery by providing insights into the quality of care across different facilities. While the integration of geospatial technology into healthcare raises privacy concerns that need addressing, its potential to optimise healthcare systems globally and contribute to improved public health outcomes is undeniable (Faruque, 2022). It is crucial for the healthcare industry to navigate the challenges and ethical considerations associated with geospatial intelligence, ensuring the responsible use of location-based data while harnessing its power to revolutionise healthcare delivery and management.
3.3.3.3 Urban Planning and Management
Geospatial technology plays a pivotal role in shaping smart cities, providing a robust foundation for enhanced urban management. Through the integration of diverse geospatial tools such as GIS, GPS, remote sensing, and 3D modeling, smart cities can seamlessly collect real-time data on various facets of their environment, infrastructure, and citizen activities. Geospatial technology facilitates a spectrum of applications, from urban planning and transportation management to emergency services and environmental monitoring. For instance, it enables the creation of dynamic maps that aid in economic development, event management, and route planning. Moreover, this technology supports initiatives like renewable energy mapping, hazard identification, and infrastructure planning, contributing to the sustainable growth and resilience of modern urban landscapes (Klei, 2019).
Balancing the benefits of smart city technologies with the imperative to protect individual privacy becomes crucial. This necessitates thoughtful considerations regarding data governance, transparency, user consent, and ethical use of surveillance technologies to prevent unwarranted intrusions into individuals’ private lives. Efforts to address these privacy challenges must involve setting balanced default parameters, providing users with control over data settings, and promoting aggregate data processing to ensure anonymity, all while maintaining the delicate equilibrium between technological innovation and individual privacy rights (Fabrègue and Bogoni, 2023).
3.3.3.4 Agriculture
Geospatial technology plays a pivotal role in agriculture. Through the utilisation of remotely sensed images acquired from satellites, aerial platforms, and GPS-tagged drones, geospatial technology facilitates the mapping and monitoring of various agricultural parameters, including soil moisture, nutrient availability, crop types, growth stages, and disease stress. The integration of geospatial technology into precision agriculture not only supports sustainable development goals but also aids in conserving biodiversity, mitigating environmental impacts, and promoting sustainable land management practices worldwide (Pandey and Pandey, 2023; Obi et al., 2023).
Privacy concerns in agriculture arise from the widespread adoption of geospatial technology and digital farming practices. The use of GPS-linked data in digital farming, aimed at achieving sustainable agriculture, poses risks of re-identification and disclosure of private information due to the spatial nature of agricultural data. To mitigate these concerns, spatial anonymisation, obfuscation, and geomasking methods are used (Nowbakht, 2022).
4. Existing literature on geomasking techniques
In this section, a comprehensive survey of recent geomasking techniques is conducted, examining the methods employed to preserve geospatial privacy while maintaining data utility and accuracy. We provide the details of the existing literature following the categorisation in Section 3.1.
4.1. Random perturbation
Table 4. Existing literature with masking using random perturbation techniques.
Density adaptive displacement Uniform population density distribution
Circular masking with fixed radius
(Kwan et al., 2004a)
Circular masking with random radius
(Armstrong et al., 1999; Kwan et al., 2004a)
Random directions within annual (Donut masking)
(Stinchcomb, 2004; Allshouse et al., 2010; Hampton et al., 2010)
Heterogeneous population density distribution
Spatially adaptive random perturbation (SARP)
(Lu et al., 2012)
Weighted random perturbation (WRP)
(Allshouse et al., 2010)
Gaussian deviation for displacement magnitude Gaussian displacement (Cassa et al., 2008; Fanshawe and Diggle, 2011; Zandbergen, 2014)
Bimodal Gaussian displacement
(Cassa et al., 2006; Zandbergen, 2014)
Density based Gaussian spatial skew
(Cassa et al., 2006)
Multiple risk factor adaptive displacement
Triangular Displacement (Murad et al., 2014)
Table 4 outlines different papers that proposed or utilised a random perturbation technique. Kwan et al. (Kwan et al., 2004a) investigate the use of circular masking with a fixed and random radius $r$. According to the investigation, a larger $r$ for the same mask leads to more deviation from the pattern obtained from the unmasked data. Further, the study suggests that circular and random radius masks, in general, yield better results than the weighted masks with the same $r$ (Allshouse et al., 2010).
Stinchcomb (Stinchcomb, 2004) was the first to explore donut masking for displacing points in random directions within an annulus. However, in cases where population density is not evenly distributed, donut masking could lead to revealing information about individuals. To mitigate this, Allshouse et al. (Allshouse et al., 2010) suggest tripling the minimum displacement area within the donut for heterogeneous populations, ensuring privacy preservation with less than a 1% error rate.
Further, Hampton et al. (Hampton et al., 2010) examine the donut method for disease mapping and compares its performance with other geomasking techniques (i.e., aggregation and random perturbation) in terms of cluster detection and patient geoprivacy protection. Their investigation shows that the donut method and random perturbation outperformed aggregation in measuring hot spot/clustered areas. However, the donut method provides a higher level of privacy preservation compared to the random perturbation, with only a slight reduction in cluster detection performance, especially in areas where individual privacy is at a higher risk.
Lu et al. (Lu et al., 2012) explore a masking technique based on Spatially adaptive random perturbation (SARP). As we have discussed in Section 3.1, the SARP technique defines the perturbation zone based on the actual distribution of residential addresses (risk location) instead of people (risk population). It employs ”donut-shaped” spatially varied perturbation zones rather than ”pancake-shaped” which allows a point to be moved to a nearby location within its immediate surroundings. The study findings show that SARP geomasking, utilising actual street addresses, offers better privacy protection compared to geomasking based on population size. While the SARP techniques do not significantly alter clustering patterns on a global scale, the geomasked data tends to exhibit greater clustering compared to the real distribution.
Cassa et al. (Cassa et al., 2006), Zandbergen (Zandbergen, 2014), and Cassa et al.(Cassa et al., 2008) investigate the masking methods based on Gaussian displacement. However, these studies highlight that the main challenge of Gaussian displacement lies in the limited point displacement due to its concentration around the mean. As highlighted by(Zandbergen, 2014), donut masking and bimodal Gaussian displacement are very similar in terms of the general area where masked locations are placed relative to the original locations.
Murad et al. (Murad et al., 2014) investigate the use of triangular displacement as a masking technique for geodata. The authors suggest a method that evaluates various risk factors to determine the appropriate displacement distances for individual location points. The method considers the sensitivity level of the data, adding specific displacement distances for low and high sensitivity information. Quasi-identifiers, which pose a combined risk when aggregated, prompt the algorithm to add additional distance to the displacement distance, while grouping these quasi-identifiers until a preset number of individuals are in each group. The algorithm then considers the intended release of the dataset, differentiating between public and restricted access. The resulting displacement distance, treated as a range with lower and upper limits ($R_{Lower}$, $R_{Upper}$), is employed in the Triangular Displacement method. This method, based on the Pythagorean Equation, introduces a random yet controlled shift in the original location points, providing a versatile and context-aware approach to spatial data masking.
4.2. Masking with pre-defined locations
Zhang et al. (Zhang et al., 2017) was the first to use location swapping as a masking technique for geospatial data in the health analytic domain. This technique offers a more realistic scenario for displacement location selection. However, as the authors highlighted, if an attacker is able to determine the buffer distance used to relocate a point, it might be possible to re-identify the original locations. The same authors also explored the location swapping with donut masking to set a minimum distance preventing points from being swapped with any locations within this specified range. The authors employed an internal buffer size half that of the external buffer.
Swanlund et al. (Swanlund et al., 2020b) introduce street masking, which uses a road network as a basis for masking. The authors use OpenStreetMap road network to replace the original location with coordinates sampled from a set within a predefined radius. As the study suggests, street masking performs similarly to population-based donut geomasking with regard to privacy protection, achieving comparable k-anonymity values at similar median displacement distances.
Geo-indistinguishability, as proposed by Andrés et al. (Andrés et al., 2013), is a privacy notion proposed for preserving privacy in location-based services. It offers stronger protection than spatial k-anonymity by safeguarding against adversaries with any current or future side information. The notion is formalised as a user enjoying $\epsilon r$-privacy within a radius $r$, where $\epsilon$ determines the level of privacy per unit distance, with smaller values of $\epsilon$ providing higher privacy protection. It achieves privacy protection by relocating individual point locations, ensuring that within a certain radius, any two points yield similar locations after relocation.
Lin (Lin, 2023) introduces geo-indistinguishable masking as a robust solution to safeguard privacy in spatial point mapping. By leveraging the concept of geo-indistinguishability, the approach ensures strong privacy protection while preserving spatial point patterns through a two-step process. Initially, masking areas are generated around each spatial point, comprising a set of candidate locations for potential relocation. Subsequently, an optimisation model is employed to strategically select the optimal candidate location within each masking area, minimising displacement distance while adhering to the stringent privacy criterion of geo-indistinguishability. While sharing similarities with location swapping methods, which also relocate spatial points within masking areas, geo-indistinguishable masking distinguishes itself by prioritising optimal candidate selection over random assignment, thereby offering superior privacy guarantees. Despite its computational complexity and potential limitations regarding temporally correlated spatial points and risk of false identification, the method represents a significant advancement in the realm of privacy-preserving spatial analysis, essential for upholding individual privacy rights in an era of burgeoning geospatial data utilisation.
4.3. Affine transformation
Recently, Wang et al. (Wang et al., 2022) conduct a study that investigates the use of translation, rotating, and scaling techniques for geomasking. The authors highlight that the coordinates of the original records can be easily recovered from the geomasked dataset if the attacker knows the transformation matrix, and even the identification of only a few points could make it possible to re-identify the entire point pattern.
Leitner and Curtis were the first to explore flipping as a geomasking technique (Leitner and Curtis, 2004). The authors state that most variants (e.g. flipping horizontally and flipping vertically) of the flipping method are designed for publishing maps and not for releasing the coordinates (Leitner and Curtis, 2006). Broen et al. (Broen et al., 2021) investigate the impact of horizontal shear as a masking technique. According to the authors, horizontal shear confers the greatest anonymity using the k-anonymity metric and maintains some spatial patterns compared to other affine transformation techniques. However, the method is not secure, as points can be trivially re-identified if the angle of the shearing can be determined.
4.4. Spatial aggregation and Clustering
Table 5 outlines the literature that uses spatial aggregation and clustering techniques for masking geodata.
Table 5. Existing literature with Spatial aggregation and clustering
Grid based aggregation Fixed size grid cells Grid line masking (Leitner and Curtis, 2004; Gupta and Rao, 2020)
Grid center masking (Leitner and Curtis, 2004; Gupta and Rao, 2020)
Varying size grid cells
Quadtree hierarchical
geographic data structure
(Lagonigro et al., 2017)
Military Grid Reference System (MGRS) (Clarke, 2016)
Spatial Tessellation based techniques Voronoi masking (Seidl et al., 2015)
Adaptive areal elimination (AAE)
(Kounadi and Leitner, 2016)
Adaptive areal masking (AAM)
(Charleux and Schofield, 2020)
Adaptive voronoi masking (AVM)
(Polzin, 2020)
RDV masking
(Gupta and Rao, 2020)
Zip4-codes aggregation (Ajayakumar et al., 2023)
Differentially private clustering (Lu and Shen, 2020; Su et al., 2017; Balcan et al., 2017; Ghazi et al., 2020; Cohen et al., 2021; Zhang et al., 2021; Jones et al., 2021; Ni et al., 2021; Chen et al., 2023)
Aggregation and perturbation (Houfaf-Khoufaf et al., 2021)
As explored by (Leitner and Curtis, 2004; Gupta and Rao, 2020), grid masking disparate every original location data point to uniform grid cells. However, grid masking disrupts the original data points’ pattern and reduces the data utility for further research and analysis. Lagonigro et al. (Lagonigro et al., 2017) propose a methodology based on quadtree hierarchical geographic data structures to create a varying-size grid adapted to local area densities. As a hierarchical grid system, it can easily be aggregated into a fixed-size grid on a lower resolution level. The methodology has been implemented using the R software and applied to the Catalonian 2014 population register data. Clarke (Clarke, 2016) investigates a multiscale geomasking technique, whereby locations are converted to Military Grid Reference System (MGRS) coordinates, providing a unique level of control over the adjusted locations. Despite the computing time required for processing, the method offers new possibilities for protecting sensitive geospatial data, overcoming disadvantages of geometric point masking methods.
Voronoi masking (VM) is a spatial tessellation technique that can retain the spatial integrity of the original data. Seidl et al. (Seidl et al., 2015) was the first to introduce the idea of VM. One advantage of this method is that points in neighbouring polygons are displaced to the same position, enhancing their K-anonymity. However, VM can dislocate some data points to large distances if the area contains scatted positions, which will change spatial patterns. This mask was automated in Python for ArcGIS Pro and ArcMap.
Kounadi and Leitner (Kounadi and Leitner, 2016) propose Adaptive areal elimination (AAE) which creates areas of a minimum K-anonymity, and then original points are either randomly perturbed within the areas or aggregated to the median centers of the areas. According to the authors, AAE introduces less spatial error than the donut mask for all the measures of spatial error. Polzin (Polzin, 2020) integrates the principles of AAE and VM to introduce a masking technique called Adaptive voronoi masking (AVM). AVM addresses topographical errors by relocating data points to the nearest street intersection, minimising false re-identification and illogical placements in water bodies or forests.
Gupta and Rao (Gupta and Rao, 2020) propose a method very similar to VM called RDV masking. Depending on the region size, for large regions, points are moved to the nearest segment of the edges of the polygon. Charleux and Schofield (Charleux and Schofield, 2020) introduce a variation of AAE called Adaptive Areal Masking (AAM). This method tries to merge polygons not by the largest shared border but with the polygon with the closest centroid. However, in configurations where the number of points to anonymise is high enough and the number of population polygons is low enough, AAE can outperform AAM. Similar to AAE, anonymisation polygons generated by the AAM method should not be disclosed to end users to minimise any potential reidentification attacks.
Ajayakumar et al. (Ajayakumar et al., 2023) propose to use Zip4 codes as a spatial aggregation unit for sharing fine-scale health data. Analysing 1.6 million Ohio Zip4 codes, the research reveals Zip4’s high spatial precision, particularly in urban locales, making it well-suited for detailed spatial analysis than common aggregation units. However, the study underscores a potential privacy risk, noting that Zip4 centroids have a higher likelihood of compromising spatial anonymity, with 73% of addresses exhibiting a low spatial k-anonymity value.
Differential private (DP) clustering focuses on the critical challenge of clustering sensitive data while preserving individual privacy, which holds significant importance in machine learning and data analysis applications. Su et al. (Su et al., 2017) delve into the limitations of current state-of-the-art approaches for k-means clustering, particularly focusing on methodologies that employ a single-workload technique. Ghazi et al. (Ghazi et al., 2020) delve into differentially private clustering, focusing on fundamental clustering problems such as Euclidean DensestBall, 1-Cluster, k-means, and k-median. Cohen et al. (Cohen et al., 2021) and Lu and Shen (Lu and Shen, 2020) propose a new approach to integrating differential privacy into clustering models, addressing the limitations of existing differentially private clustering algorithms. Yang et al. (Yang et al., 2022) address the emerging field of privacy-preserving data analysis, specifically focusing on constructing a privacy-preserving K-means clustering protocol that ensures local privacy for user data.
Houfaf-Khoufaf et al. (Houfaf-Khoufaf et al., 2021) study the combination of aggregation techniques with perturbation mechanisms. The study also introduces the simulated cluster crowding (SCC) method, a novel approach inspired by the simulated crowding of bike GPS tracks. SCC generates new random points within existing clusters’ extents to achieve both high k-anonymity and cluster preservation. The clusters are initially generated using the DBSCAN algorithm, distinguishing between outlier points and those within clusters. Outlier points are masked using the bimodal Gaussian perturbation method, while points within clusters are masked by generating the same number of new points randomly inside the polygon extent of each cluster. Three versions of SCC are proposed: SCCv1 generates points uniformly within the cluster area, SCCv2 introduces buffer areas around initial points to ensure a minimum distance, and SCCv3 dilates roads and intersects them with clusters to generate fake points. The methods are evaluated based on k-anonymity, cluster preservation, and privacy guarantee, demonstrating their effectiveness in maintaining privacy while preserving the spatial distribution of clusters.
4.5. Synthetic data generation
Quick and Waller (Quick and Waller, 2018) propose the use of spatiotemporal data analysis methods for generating synthetic data to address disclosure risks associated with releasing public-use data in disease mapping, where small counts present challenges. The study employs a Bayesian model where the synthetic data are generated from the resulting posterior predictive distribution, preserving different dependencies. Drechsler and Hu (Drechsler and Hu, 2021) investigate the viability of using synthetic data generation to provide detailed geocoding information from a large administrative database while maintaining confidentiality. The authors show that a synthesis strategy based on categorical CART models performs well in addressing the risk-utility trade-off.
Quick (Quick, 2021) focuses on the dissemination of synthetic data as an effective means of making sensitive information publicly available while minimising the risk of disclosure. It extends a differential privacy approach to Poisson-distributed count data, accommodating population size heterogeneity and incorporating prior information on event rates. Hong et al. (Hong et al., 2022) propose a privacy-preserving synthetic data generation framework for addressing privacy concerns associated with sharing complete datasets. The method involves synthesising a new privacy-preserving dataset through a generative model that captures the probability density function of the original dataset’s features. The synthetic dataset is designed to be statistically distinct from the original while preserving data utility.
Koebe et al. (Koebe et al., 2023) propose an innovative microdata dissemination strategy for household survey data to balance privacy and data utility. Instead of traditional approaches involving perturbed cluster locations, they suggest publishing two datasets: the first contains original survey data with limited geographic identifiers, suitable for representative-level analysis; the second involves synthetically generated survey data with accurate cluster locations, enhancing proximity-related information for spatial analysis. The proposed strategy significantly reduces re-identification risks compared to common spatial perturbation methods, while maintaining data utility for both non-spatial and spatial analyses. The authors employ copulas for synthetic data generation, emphasising their computational efficiency, interpretability, and practicality for data producers. The methodology is seen as beneficial for expanding the use cases of survey data products, supporting mapping initiatives, and ensuring data access while complying with privacy regulations.
5. Open issues and future directions
As we have reviewed in Section 4, no geomasking technique is complete and capable of providing high accuracy while ensuring strong privacy against different attacks. Each masking technique has distinct advantages and disadvantages, requiring data custodians to understand these techniques and their capabilities before applying them in practical applications. While most masking literature provides a detailed explanation of how to use the masking procedure in general, most masking methods lack a corresponding guide for the optimal choice of parameters, as these parameter settings need to be changed according to the data that needs protection. Further, some papers do not take into account different scenarios that could occur when their proposed masking techniques are applied in real applications. For example, a technique such as location swapping could lead to issues when no residential addresses are found within the radius that is defined based on the population density of a region (Fabrègue and Bogoni, 2023).
The main question in geomasking research is how to balance the need for anonymisation with the usefulness of geodata while also considering the probability and potential harms of re-identification (Helderop et al., 2023). In general, geomasking methods can preserve data accuracy but are presumed to impose a high risk of re-identification. As discussed in Section 3.1, complete anonymisation is impossible if the data is to remain useful. Therefore, research efforts should focus not only on better preventing identification but also on considering the probability of re-identification events and the potential harms of such incidents.
False identification is an emerging concern of applying geomasking methods, which indicate the linking of the masked data points to incorrect persons or households (Seidl et al., 2018). The false identification transferred the potential negative effects of being identified from the true persons or households to individuals who were not part of the research. Such limitations influence both the disclosure risk and a successful investigation of spatial patterns. There are newly developed geomasking methods that target addressing this issue, such as adaptive Voronoi masking (Kounadi and Leitner, 2016). Though these techniques might be better approaches to visualising a protected version of the distribution of a point pattern, however, they are less accurate in detecting local patterns in location data. Further, these techniques might not be applicable to different geodata types that are complex, such as spatiotemporal stamps of a user in a social network.
As we have discussed in Section 3.1, many geomasking techniques could not provide provable privacy guarantees as their underlying methodologies only contain non-provable privacy techniques, such as k-anonymity (Wang et al., 2022; Broen et al., 2021). Techniques such as differential privacy have been used in location-based services (Yan et al., 2022) for providing strong and provable privacy guarantees when answering queries, but have limited adoption and utilisation in the masking of geodata (Harris, 2020). For example, to improve geodata anonymisation and align with the relevant risk factors, it may have been more appropriate to add random noise to the geographic information of routes mapped by just a few individuals, or exclude routes mapped by very few individuals (e.g., publish the number of routes in the area, but not the exact paths).
In the geomasking literature, each paper tends to use different datasets and different measures to evaluate their proposed methodologies. As shown by Redlich (Redlich, 2022), the choice of different accuracy and privacy metrics influence the evaluation of the geomasking methods, leading to different conclusions. The utilisation of different experimental settings in geomasking literature could potentially make the comparison of different geomasking techniques difficult. Thus, having a benchmark setting, including a set of datasets, accuracy measures, and privacy measures, would allow for more accurate quantification of the performance of different geomasking methods.
Table 6. Tools designed to integrate large language models with geospatial data.
Tool Description
IBM Geospatial foundation model  (Raghavan and Shim, 2023) In collaboration with NASA, IBM has developed the first-ever foundation model for analysing geospatial data, watsonx.ai model, to convert satellite data into high-resolution maps. The model aims to assist in estimating climate-related risks, monitoring forests for carbon-offset programs, and creating predictive models for climate change adaptation. This model, based on NASA’s Harmonised Landsat Sentinel-2 data, is trained to understand satellite images, recognising features like historic floods and fire burn scars.
ChatGeoPT (Strong, 2023) ChatGeoPT, a Geospatial AI Assistant, is a prototype that explores the integration of LLMs into geospatial data science workflows. The application accepts natural language prompts and translates them into queries for the OpenStreetMap (OSM). ChatGeoPT is seen as a preliminary step toward the development of a more extensive Large Earth Observation Model (LEOM), aiming to enable intelligent interactions and summarisations of geospatial data from diverse sources, including satellite imagery and weather data, extending beyond the capabilities of OpenStreetMap.
MapsGPT (Proxi, nd) MapsGPT simplifies the process of creating interactive maps, eliminating the need for extensive coding skills and making it accessible to everyone. Unlike traditional GIS, MapsGPT allows users to generate customised maps effortlessly by inputting a simple sentence. MapsGPT provides a versatile solution for discovering and curating locations, making it easy to plan activities, find essential stores, and explore a new city.
LLM-Geo (Li and Ning, 2023) LLM-Geo represents an innovative paradigm known as Autonomous GIS, introducing an AI-powered geographic information system that leverages LLMs like ChatGPT to transform spatial analysis. LLM-Geo seeks to automate spatial data collection, analysis, and visualisation by harnessing the natural language understanding, reasoning, and coding capabilities of LLMs.
GEoLM (Li et al., 2023) GEoLM, a geospatially grounded language model, enhances natural language understanding by integrating valuable geospatial information from extensive databases like OSM. GEoLM connects linguistic details in textual corpora with geospatial data retrieved from geographical databases. GEOLM stands as a versatile tool for various geographically related language understanding tasks, with future plans to extend its applications to question-answering and auto-regressive tasks.
GFM (Mendieta et al., 2023) GFM develops geospatial foundation models for applications like agriculture, urban planning, and disaster response. The authors introduce GeoPile, a compact and diverse dataset sourced from various remote sensing platforms, to enhance feature diversity during pretraining. They explore the potential of continual pretraining from readily available ImageNet-22k models.
5.1. Integration of large language models (LLMs) with Geospatial data
In recent years, substantial progress has been achieved in the development of foundational models. Notably, the integration of these foundational models with geospatial data, referred to as GeoAI Foundation Models, has demonstrated significant potential across various applications, ranging from geographic question answering and remote sensing image understanding to map generation and location-based services (Zhang et al., 2023b). Nevertheless, it is crucial to acknowledge that the development and application of GeoAI foundation models introduce significant privacy and security risks, an aspect that has not yet received comprehensive attention to date.
Rao et al.(Rao et al., 2023) examine the privacy and security risks associated with the entire lifecycle of GeoAI foundation models. The study identifies a range of potential privacy and security risks in different phases of GeoAI foundation model development. The authors highlight that leveraging such foundational models in geospatial research poses challenges, particularly in balancing privacy and utility, mitigating geographic biases, and addressing cross-modal privacy risks.
Several other studies have explored the use of foundation models in geospatial data analysis. Mooney et al. (Mooney et al., 2023) investigate the performance of ChatGPT in a real Geographic Information Systems (GIS) exam. The study identifies challenges in teaching spatial concepts to LLMs, mainly due to the limitations in handling multimodal content. Tao and Xu (Tao and Xu, 2023) explore the potential of leveraging ChatGPT for the task of map design. The authors conduct two sets of experiments, testing the generation of thematic maps through specified prompts and the creation of mental maps purely from textual descriptions of geographic spaces. Though ChatGPT’s reliance on textual input and output, the authors show that ChatGPT holds the potential to revolutionise map design processes, particularly in educational and research contexts, while emphasising the need for cautious consideration of its current limitations and validation of data sources.
Similarly, Zhang et al. (Zhang et al., 2023b) propose GeoGPT, a framework that integrates the semantic understanding capabilities of LLMs. GeoGPT utilises the Langchain framework to connect the LLM with a pool of GIS tools, allowing it to choose and execute appropriate tools for tasks like geospatial data collection, spatial analysis, and mapping. Zhang et al. (Zhang et al., 2023a) explore the ethical dimensions of leveraging generative AI in the domain of cartography. Focusing specifically on maps generated by LLMs, the researchers establish an open-source dataset containing both AI-generated and human-designed maps. The authors show that the maps generated by the LLM lack geographic coverage and diversity. However, they highlight that such AI tools could be valuable in research settings where real geodata is limited. Table 6 highlights some other tools specifically designed to integrate LLMs with geospatial data.
6. Conclusion
In this paper, we present a survey of historical and current state-of-the-art masking techniques for geodata. We proposed a taxonomy that serves as a comparison and analysis tool for geomasking techniques. In this taxonomy we identified three main dimensions: privacy techniques, evaluation measures, and practical aspects. These dimensions enabled us to characterize geomasking techniques and generate a taxonomy of such techniques. Through this taxonomy, we identified various shortcomings in current approaches to geomasking, suggesting several future research directions in this field. An area for future research involves investigating how large language models (LLMs) can be employed to analyse and enhance geomasking techniques. To date, no studies have delved into the potential integration of LLMs for this purpose, making it a promising area for future exploration and improvement. Solving these open research questions is a core requirement to make geomasking applicable for practical applications.